# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 8. KSBA moduli spaces of stable log Calabi–Yau surfaces
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface and let $\mathscr{P}$ be a good polyhedral decomposition of a Symington polytope $P$ associated to $(Y,D,L)$ as in §2.3. Let $\pi_{\mathscr{P}}\colon\mathcal{X}_{\mathscr{P}}\to\Delta$ be the semistable mirror to the maximal degeneration of $(Y,D,L)$ defined by $\mathscr{P}$, and denote by
$(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\longrightarrow S_{\mathcal{X}_{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{\overline{X}}_{\mathscr{P}})]$
the polarized mirror family as in Definition 6.17. In what follows we refer to this family as the “double mirror” to $(Y,D,L)$. In this section we first investigate in Theorem 8.2 how the double mirror changes if $(P,\mathscr{P})$ is rescaled. We then use this result to show in Theorem 8.6 that, up to rescaling $(P,\mathscr{P})$, the restriction of the double mirror to a particular one-dimensional locus is a toric degeneration. Using this, we show in Theorem 8.10 that the general fiber of the double mirror is deformation equivalent to $(Y,D,L)$. Consequently, we prove in Theorem 8.11 that we obtain a surjective and finite map from the toric variety $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ to the coarse KSBA moduli space $M_{(Y,D,L)}$.
@@ 8.1. Base change of open Kulikov degenerations
Note that $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is not an open Kulikov degeneration in general, as the total space $\mathcal{X}_{k\mathscr{P}}$ is singular. Instead, $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is log smooth for the divisorial log structure on $\mathcal{X}_{k\mathscr{P}}$ defined by the central fiber $\mathcal{X}_{k\mathscr{P},0}:=\pi_{k\mathscr{P}}^{-1}(0)$, and the divisorial log structure on $\Delta$ defined by the origin $0\in\Delta$. Moreover, the dual intersection complex of the central fiber $\mathcal{X}_{k\mathscr{P},0}$ is the scaled polyhedral decomposition $k\mathscr{P}$ on the scaled polytope $kP$. The following result will allow us to construct an open Kulikov degeneration which is a resolution of singularities of $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$.
@@
We successively pull, as in the proof of Theorem 2.14, all integral points in $kP$ which are not vertices of $k\mathscr{P}$. This gives us a desired good polyhedral decomposition $\widetilde{\mathscr{P}}$. ∎
@@
$\nu_{\star}:NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\longrightarrow NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$
be the composition of the pushforward of curve classes along $\nu$ with the natural isomorphism $NE(\mathcal{X}_{k\mathscr{P}}/\overline{\mathcal{X}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$ whose existence follows from the fact that the base change $t\mapsto t^{k}$ does not change the central fiber, and so $\mathcal{X}_{k\mathscr{P},0}\simeq\mathcal{X}_{\mathscr{P},0}$. We denote by $\rho_{\nu}$ the scheme morphism
$\rho_{\nu}:\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})]\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})]$
@@
$(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{\overline{X}}_{\mathscr{P}})]$
@@
$\rho_{\nu}:\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})]\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})]$
@@
$(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{\overline{X}}_{\widetilde{\mathscr{P}}})]$
@@
It suffices to identify the theta bases and the structure constants of the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the generators and structure constants of the subalgebra of the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ consisting of elements of degree divisible by $k$. To identify the theta bases, one notes that the index set of the theta basis for the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is $C(kP)_{\mathbb{Z}}$, which is exactly the set of integral points of height divisible by $k$ in the cone $C(P)_{\mathbb{Z}}$ indexing the theta basis for the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Hence, it remains to show that, for every $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$, we have
$N_{pqr}^{\beta}(\mathcal{X}_{\mathscr{P}})=\sum_{\begin{subarray}{c}\beta^{\prime}\in NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\\ \nu_{\star}\beta^{\prime}=\beta\end{subarray}}N_{pqr}^{\beta^{\prime}}(\mathcal{X}_{\widetilde{\mathscr{P}}})\,.$ (8.1)
To prove Equation (8.1), we first note that, even if the base change $\mathcal{X}_{k\mathscr{P}}$ of $\mathcal{X}_{\mathscr{P}}$ is not the total space of an open Kulikov degeneration, it is still log smooth over $\Delta$, and so one can also define invariants $N_{pqr}^{\beta}(\mathcal{X}_{k\mathscr{P}})$ for $\mathcal{X}_{k\mathscr{P}}$. As the base change $t\mapsto t^{k}$ does not change the central fiber and only scales the log structure, it follows that we have
@@
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\overline{\mathcal{X}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\overline{\mathcal{X}}_{\mathscr{P}})$. On the other hand, as $\nu:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{k\mathscr{P}}$ is a log modification, it follows from [56, §10] that
$N_{pqr}^{\beta}(\mathcal{X}_{k\mathscr{P}})=\sum_{\begin{subarray}{c}\beta^{\prime}\in NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\overline{\mathcal{X}}_{\widetilde{\mathscr{P}}})\\ \nu_{\star}\beta^{\prime}=\beta\end{subarray}}N_{pqr}^{\beta^{\prime}}(\mathcal{X}_{\widetilde{\mathscr{P}}})\,,$ (8.3)
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\overline{\mathcal{X}}_{k\mathscr{P}})$. Since Equation (8.1) follows from Equations (8.2) and (8.3), this concludes the proof. ∎
@@ 8.2. From semistable mirrors to toric degenerations
Let $(Y,D,L)$ be a polarized log Calabi–Yau surface with a maximal degeneration defined by $(P,\mathscr{P})$ as in §2.4 and let $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\to\Delta$ be the corresponding semistable mirror as in §4.3.
We use the notation introduced in the construction of $(P,\mathscr{P})$ in §2.2. For every $1\leq i\leq n$ and $1\leq j\leq N_{i}$, consider the image $L_{ij}$ in $P$ of the line in $\overline{P}$ passing through $p_{ij}$, and parallel to the edge $\overline{P}_{i}$ of $\overline{P}$. By Lemma 2.11, the intersection of $L_{ij}$ with the interior of $P$ is an interval of positive length, and so one can find a rational point $p_{ij}^{\prime}\in L_{ij}$ close enough to $p_{ij}$ in the sense that the smallest cell of $\mathscr{P}$ containing $p_{ij}^{\prime}$ also contains $p_{ij}$, and such that all the points $p_{ij}^{\prime}$ are distinct and contained in the interior of $P$.
Fix a large enough integer $k$ such that all the points $p_{ij}^{\prime}$ become integral points of the scaled polytope $kP$. Arguing as in the proof of Lemma 8.1 and pulling first the points $p_{ij}^{\prime}$ among the integral points of $kP$ which are not vertices of $k\mathscr{P}$, we obtain a good polyhedral decomposition $\widetilde{\mathscr{P}}$ on $kP$, obtained as a refinement of $k\mathscr{P}$, and such that every line segment $[p_{ij},p_{ij}^{\prime}]$ is a union of edges of $\widetilde{\mathscr{P}}$.
@@
Figure 8.1. On the left, triangles $\Delta_{ij}$ cut out from $\overline{P}$. In the center, induced cuts in $P$. On the right, deformed cuts $\delta_{ij}$.
Continuously moving the tips of the cut triangles $\Delta_{ij}$ on the line $L_{ij}$ away from $p_{ij}$ towards $p_{ij}^{\prime}$ produces a deformation of the integral affine structure on $P$, with endpoint an integral affine structure with distinct focus-focus singularities $\widetilde{p}_{ij}$ – see Figure 8.1. Moreover, we can assume that $\widetilde{p}_{ij}$ is contained in the interior of the edge of $\widetilde{\mathscr{P}}$ containing $p_{ij}$ and contained in $[p_{ij},p_{ij}^{\prime}]$. In particular, the cuts induced by the triangles $\Delta_{ij}$ deform into lines $\delta_{ij}$ connecting the side $\overline{P}_{i}$ to the focus-focus singularities $\widetilde{p}_{ij}$. Every integral point $v\in P_{\mathbb{Z}}$ is smooth for the deformed integral affine structure, and so its star defines a toric fan $\Sigma_{v}$. Note that, for every $i,j$ such that $v=p_{ij}$, the line segment $[p_{ij},p_{ij}^{\prime}]$ corresponds to a ray $\rho_{ij}$ of the fan $\Sigma_{v}$.
@@
Recall from §4.2 that the log Calabi–Yau surface $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is a smoothing of the toric surface $(X^{v}_{n\mathscr{P}},\partial X^{v}_{n\mathscr{P}})$ with fan the star of $v$ in $\overline{P}$. Using the almost toric description of smoothings of quotient singularities – see [91, §6.1] and [32, Remark 2.7], it follows that $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ admits an almost toric fibration over a Symington polytope obtained from a toric polytope of $(X^{v}_{k\overline{\mathscr{P}}},\partial X^{v}_{k\overline{\mathscr{P}}})$ by node smoothing. Therefore, by the combinatorial duality between internal blow-ups and node smoothings – see e.g.  [30, p 488], the pair $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is diffeomorphic, and so deformation-equivalent by [34, Theorem 5.14], to the log Calabi–Yau surface described in Lemma 8.3 as internal blow-ups of the toric surface $(X_{\Sigma_{v}},\partial X_{\Sigma_{v}})$. ∎
We denote by $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ the normal crossing surface obtained from $\mathcal{X}_{\widetilde{\mathscr{P}},0}$ by blowing down all the $(-1)$-curves as in Lemma 8.3. We show below that $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is a central fiber of a toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ in the sense of [47], that is, with toric irreducible components of the central fiber glued along toric divisors, and with the map to $\Delta$tale locally given by a toric morphism near the 0-dimensional strata of the central fiber.
@@
The blowdown morphism $p_{0}:\mathcal{X}_{\widetilde{\mathscr{P}},0}\to\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is projective and deforms to a projective birational morphism $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\to\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$, contracting finitely many internal exceptional curves of the central fiber, and fitting into a commutative diagram
[…]
@@
Since $\widetilde{\mathscr{P}}$ is a good polyhedral decomposition on $kP$ as in §2.3, there is a strictly convex integral multi-valued PL function $\varphi$ on $kP$ whose domains of linearity agree with $\widetilde{\mathscr{P}}$, as in §2.4. Then, we obtain from $\varphi$ a strictly convex integral multi-valued PL function $\varphi^{gs}$ on $kP$ on the integral affine manifold with singularities $(kP,\widetilde{\mathscr{P}})^{gs}$ obtained from $(kP,\widetilde{\mathscr{P}})$ by pushing singularities away from the vertices as described before Lemma 8.3 – see [18, §3] where such PL functions on integral affine manifolds with singularities pushed away are studied in detail. Therefore, by [47, Theorem 2.34] since $P$ is a disk with second cohomology zero, $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is a quasi-projective scheme. Hence, we have an ample line bundle $L_{0}$ on $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ whose pullback under $\pi_{\widetilde{\mathscr{P}}}$ is a nef line bundle on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $H^{2}(\mathcal{X}_{\widetilde{\mathscr{P}},0},\mathcal{O}_{\mathcal{X}_{\widetilde{\mathscr{P}},0}})=0$ by Lemma 3.10, this line bundle deforms to a nef line bundle $L$ on $\mathcal{X}_{\widetilde{\mathscr{P}}}$. Since $\mathcal{X}_{\widetilde{\mathscr{P}}}$ is a Mori dream space by Proposition 3.36, there is a projective birational morphism $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ exactly contracting the curves $C$ with $L\cdot C=0$, that is the finitely many internal exceptional curves in the central fiber contracted by $p_{0}$.
By construction, the irreducible components of the central fiber $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ are the toric varieties $X_{\Sigma_{v}}$ as in Lemma 8.3 and are glued along toric divisors. Moreover, $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ is isomorphic to $\mathcal{X}_{\widetilde{\mathscr{P}}}$ away from the finitely many ordinary double points where the internal exceptional curves are contracted, and so in particular in the neighborhood of 0-dimensional strata of the central fiber. As $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is log smooth, it follows that $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is also log smooth near the 0-dimensional strata of the central fiber, and we conclude that $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is a toric degeneration in the sense of [47]. ∎

### v3

@@ 8. KSBA moduli spaces of stable log Calabi–Yau surfaces
Let $(Y,D,L)$ be a generic polarized log Calabi–Yau surface and let $\mathscr{P}$ be a good polyhedral decomposition of a Symington polytope $P$ associated to $(Y,D,L)$ as in §2.3. Let $\pi_{\mathscr{P}}\colon\mathcal{X}_{\mathscr{P}}\to\Delta$ be the semistable mirror to the maximal degeneration of $(Y,D,L)$ defined by $\mathscr{P}$, and denote by
$(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}},\mathcal{L}_{\mathcal{X}_{\mathscr{P}}})\longrightarrow S_{\mathcal{X}_{\mathscr{P}}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})]$
the polarized mirror family as in Definition 6.25. In what follows we refer to this family as the “double mirror” to $(Y,D,L)$. In this section we first investigate in Theorem 8.2 how the double mirror changes if $(P,\mathscr{P})$ is rescaled. We then apply this result to show in Theorem 8.7 that, up to rescaling $(P,\mathscr{P})$, the restriction of the double mirror to a particular one-dimensional locus is a toric degeneration. Building on this, Theorem 8.11 establishes that the general fiber of the double mirror is deformation equivalent to $(Y,D,L)$. Consequently, we prove in Theorem 8.13 that we obtain a surjective and finite map from the toric variety $S^{\mathrm{sec}}_{\mathcal{X}_{\mathscr{P}}}$ to the coarse KSBA moduli space $M_{(Y,D,L)}$.
@@ 8.1. Base change of open Kulikov degenerations
Note that $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is not an open Kulikov degeneration in general, as the total space $\mathcal{X}_{k\mathscr{P}}$ is singular. Instead, $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$ is log smooth for the divisorial log structure on $\mathcal{X}_{k\mathscr{P}}$ induced by the central fiber $\mathcal{X}_{k\mathscr{P},0}:=\pi_{k\mathscr{P}}^{-1}(0)$, and the divisorial log structure on $\Delta$ induced by the origin $0\in\Delta$. Moreover, the dual intersection complex of the central fiber $\mathcal{X}_{k\mathscr{P},0}$ is the scaled polyhedral decomposition $k\mathscr{P}$ on the scaled polytope $kP$. The following result will allow us to construct an open Kulikov degeneration which is a resolution of singularities of $\pi_{k\mathscr{P}}:\mathcal{X}_{k\mathscr{P}}\rightarrow\Delta$.
@@
As in the proof of Theorem 2.16, we successively pull all integral points in $kP$ which are not vertices of $k\mathscr{P}$. This gives us a desired good polyhedral decomposition $\widetilde{\mathscr{P}}$. ∎
@@
$\nu_{\star}:NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})\longrightarrow NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$
be the composition of the pushforward of curve classes along $\nu$ with the natural isomorphism $NE(\mathcal{X}_{k\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$ whose existence follows from the fact that the base change $t\mapsto t^{k}$ does not change the central fiber, and so $\mathcal{X}_{k\mathscr{P},0}\simeq\mathcal{X}_{\mathscr{P},0}$. We denote by $\rho_{\nu}$ the scheme morphism
$\rho_{\nu}:\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})]\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})]$
@@
$(\mathcal{Y}_{\mathcal{X}_{\mathscr{P}}},\mathcal{D}_{\mathcal{X}_{\mathscr{P}}})\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})]$
@@
$\rho_{\nu}:\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})]\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})]$
@@
$(\mathcal{Y}_{\mathcal{X}_{\widetilde{\mathscr{P}}}},\mathcal{D}_{\mathcal{X}_{\widetilde{\mathscr{P}}}})\longrightarrow\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}_{\widetilde{\mathscr{P}}}/\mathcal{X}^{\mathrm{can}}_{\widetilde{\mathscr{P}}})]$
@@
It suffices to identify the theta bases and the structure constants of the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ with the generators and structure constants of the subalgebra of the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ consisting of elements of degree divisible by $k$. To identify the theta bases, one notes that the index set of the theta basis for the mirror algebra of $\pi_{\widetilde{\mathscr{P}}}:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is $C(kP)_{\mathbb{Z}}$, which is exactly the set of integral points of height divisible by $k$ in the cone $C(P)_{\mathbb{Z}}$ indexing the theta basis for the mirror algebra of $\pi_{\mathscr{P}}:\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. Hence, it remains to show that, for every $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$, we have
[…]
To prove Equation (8.1), we first note that, even if the base change $\mathcal{X}_{k\mathscr{P}}$ of $\mathcal{X}_{\mathscr{P}}$ is not the total space of an open Kulikov degeneration, it is still log smooth over $\Delta$, and so one can also define invariants $N_{pqr}^{\beta}(\mathcal{X}_{k\mathscr{P}})$ for $\mathcal{X}_{k\mathscr{P}}$. The base change $t\mapsto t^{k}$ does not change the central fiber and only scales the log structure. Thus, by [62, Corollary 6.7], we obtain
@@
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{k\mathscr{P}})\simeq NE(\mathcal{X}_{\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{\mathscr{P}})$. On the other hand, as $\nu:\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\mathcal{X}_{k\mathscr{P}}$ is a log modification, it follows from [61, §10] that
[…]
for all $p,q,r\in C(kP)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}_{k\mathscr{P}}/\mathcal{X}^{\mathrm{can}}_{k\mathscr{P}})$. Since Equation (8.1) follows from Equations (8.2) and (8.3), this concludes the proof. ∎
@@ 8.2. From semistable mirrors to toric degenerations
Construction 8.3.
We use the notation introduced in the construction of $(P,\mathscr{P})$ in §2.2, by cutting off the triangles $\Delta_{ij}$ from the polygon $(\overline{P},\overline{\mathscr{P}})$. As in [33, Lemma 5.6], the tips of the triangle $\Delta_{ij}$ can be moved along the monodromy invariant directions $L_{ij}$ to split the integral affine singularities of $P$ into distinct focus-focus singularities $p_{ij}^{\prime}$ at rational points of $P$ – see Figure 10. We denote by $\Delta_{ij}^{\prime}$ the new triangles with tips $p_{ij}^{\prime}$, and by $P^{\prime}$ the resulting integral affine manifold. Let $k\in\mathbb{Z}_{\geq 0}$ be large enough such that the points $p_{ij}^{\prime}$ become integral in the scaled polytope $kP^{\prime}$. Then, $kP^{\prime}$ is obtained from $k\overline{P}$ but cutting the integral triangles $\Delta_{ij}^{\prime}$. Hence, by Lemma 8.1, there exists a good polyhedral decomposition $\widetilde{\mathscr{P}}$ on $kP$, obtained as a refinement of $k\mathscr{P}$. Up to increasing $k$ and refining $\widetilde{\mathscr{P}}$ further, we can also assume that the monodromy invariant directions $L_{ij}^{\prime}$ are unions of edges of $\widetilde{\mathscr{P}}$, and that no edge of $\widetilde{\mathscr{P}}$ connects two focus-focus singularities $p_{ij}^{\prime}$. By Lemma 2.12, for every $i,j$, there exists an edge $e_{ij}$ of $\widetilde{P}$ adjacent to $p_{ij}^{\prime}$ and contained in the monodromy invariant direction $L_{ij}^{\prime}$. Finally, we define $\widetilde{P}$ as the integral affine manifold with singularities obtained from $P$ by continuously moving the focus-focus singularities from the integral points $p_{ij}^{\prime}$ to points $\widetilde{p}_{ij}$ in the interior of the edges $e_{ij}$.
By construction, singularities of $\widetilde{P}$ are all focus-focus, are all contained in the interior of edges of $\widetilde{\mathscr{P}}$, and every edge of $\widetilde{\mathscr{P}}$ contains at most one focus-focus singularity. Every integral point $v\in\widetilde{P}$ is smooth for the integral affine structure, and so the star of $\widetilde{\mathscr{P}}$ at $v$ defines a toric fan $\Sigma_{v}$. For every $i,j$ such that $v=p_{ij}$, we denote by $\rho_{ij}$ the ray of the fan $\Sigma_{v}$ corresponding to the edge of $\widetilde{P}$ adjacent to $v$ and contained in the line segment connecting $p_{ij}$ to $\widetilde{p}_{ij}$.
Figure 10. On the left, triangles $\Delta_{ij}$ cut out from $\overline{P}$. In the center, induced cuts in $P$. On the right, deformed cuts $\delta_{ij}$.
@@
By §4.2, the log Calabi–Yau surface $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is a corner blow-up of a smoothing of the toric surface with fan the star of $v$ in $\overline{P}$. Using the almost toric description of smoothings of quotient singularities – see [95, §6.1] and [34, Remark 2.7], we deduce that $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ admits an almost toric fibration over a Symington polytope obtained from a toric polytope by node smoothing. Therefore, by the combinatorial duality between interior blow-ups and node smoothings – see e.g.  [32, p 488], the pair $(X^{v}_{\widetilde{\mathscr{P}}},\partial X^{v}_{\widetilde{\mathscr{P}}})$ is diffeomorphic, and so deformation-equivalent by [36, Theorem 5.14], to the log Calabi–Yau surface described in Lemma 8.4 as interior blow-ups of the toric surface $(X_{\Sigma_{v}},\partial X_{\Sigma_{v}})$. ∎
We denote by $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ the normal crossing surface obtained from $\mathcal{X}_{\widetilde{\mathscr{P}},0}$ by blowing down all the (1)(-1)-curves appearing as exceptional curves of the blow-ups described in Lemma 8.4. We show below that $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is a central fiber of a toric degeneration $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ in the sense of [50], that is, with toric irreducible components of the central fiber glued along toric divisors, and with the map to $\Delta$ étale locally given by a toric morphism near the 0-dimensional strata of the central fiber.
@@
The blowdown morphism $p_{0}:\mathcal{X}_{\widetilde{\mathscr{P}},0}\to\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ is projective and deforms to a projective birational morphism $p:\mathcal{X}_{\widetilde{\mathscr{P}}}\to\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$, contracting finitely many interior exceptional curves of the central fiber, and fitting into a commutative diagram
$\mathcal{X}_{\widetilde{\mathscr{P}}}$$\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$$\Delta$$\Delta$$\scriptstyle{\pi_{\widetilde{\mathscr{P}}}}$$\scriptstyle{p}$$\scriptstyle{\pi_{\widetilde{\mathscr{P}}}^{gs}}\scriptstyle{\simeq}$
@@
[…]
By construction, the irreducible components of the central fiber $\mathcal{X}_{\widetilde{\mathscr{P}},0}^{gs}$ are the toric varieties $X_{\Sigma_{v}}$ as in Lemma 8.4 and are glued along toric divisors. Moreover, $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}$ is isomorphic to $\mathcal{X}_{\widetilde{\mathscr{P}}}$ away from the finitely many ordinary double points where the interior exceptional curves are contracted, and so in particular in the neighborhood of 0-dimensional strata of the central fiber. Since $\mathcal{X}_{\widetilde{\mathscr{P}}}\rightarrow\Delta$ is log smooth, we obtain that $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is also log smooth near the 0-dimensional strata of the central fiber. Therefore, $\mathcal{X}_{\widetilde{\mathscr{P}}}^{gs}\rightarrow\Delta$ is a toric degeneration in the sense of [50]. ∎